1/2x+1/3x+5+90=180 whats the answer?
1
Expert's answer
2012-04-27T07:32:12-0400
1/(2x)+1/(3x+5)+90=180
=> x < >0, x< > -5/3
1/(2x)+1/(3x+5)=90
(3x+5+2x)/(2x)/(3x+5)=90
(5x+5)/(2x)/(3x+5)=90
(x+1)/(2x)/(3x+5)=18
(x+1)/(3x^2+5x)=36
x+1=108x^2+180x
108x^2+179x-1=0
x1,2=(-179=-sqrt(179^2+4*180))/2/108 - answer